OpenStudy (anonymous):

What is the slope and y-intercept of the graph of -x + 2y = 6?

OpenStudy (alexandervonhumboldt2):

first convert to spole-intercept form

OpenStudy (anonymous):

make everything equal to y \[2y = 6+x\] divide by 2\[y = 3 + \frac{ x }{ 2 }\] reorder terms\[y = \frac{ x }{ 2 } + 3\]

OpenStudy (alexandervonhumboldt2):

to do this add x to both sides first. then divide botth sides by 2. you will get form of y=ax+b where a is slope and b is y-intercept

OpenStudy (anonymous):



OpenStudy (alexandervonhumboldt2):

yeas

OpenStudy (anonymous):

technically, \[\frac{ x }{ 2 } = \frac{ 1 }{ 2 }x\]

sammixboo (sammixboo):

\(\color{blue}{\text{Originally Posted by}}\) @anonymark make everything equal to y \[2y = 6+x\] divide by 2\[y = 3 + \frac{ x }{ 2 }\] reorder terms\[y = \frac{ x }{ 2 } + 3\] \(\color{blue}{\text{End of Quote}}\) Also known as changing it into slope-intercept form

OpenStudy (anonymous):

what do I divide? the 2x and the x? or with the 6?

OpenStudy (anonymous):

so ... it would be 1/2; 3 ?

OpenStudy (alexandervonhumboldt2):

no first you add x to both sides: -x+2y+x=6+x 2y=6+x now divide both sides by 2 you get y=1/2x+3

OpenStudy (alexandervonhumboldt2):

y=ax+b y=1/2x+3 a is slope and b is y-intecept

OpenStudy (alexandervonhumboldt2):